### v2

# The Invisible Leash? Why RLVR May or May Not Escape Its Origin

Abstract
Recent advances in large reasoning models highlight Reinforcement Learning with Verifiable Rewards (RLVR) as a promising method for enhancing AI’s capabilities, particularly in solving complex logical tasks. However, it remains unclear whether the current practice of RLVR truly expands a model’s reasoning boundary or mainly amplifies high-reward outputs that the base model already knows for improved precision. This study presents an empirical investigation that provides fresh insights into the potential limits of the common practice of RLVR. We examine how, under current training conditions, RLVR can operate as a support-constrained optimization mechanism that may restrict the discovery of entirely original solutions, remaining constrained by the base model’s initial distribution. We also identify an entropy–reward trade-off: while the current RLVR recipe reliably enhances precision, it may progressively narrow exploration and potentially overlook correct yet underrepresented solutions. Extensive empirical experiments validate that while the current RLVR recipe consistently improves pass@1, the shrinkage of empirical support generally outweighs the expansion of empirical support under larger sampling budgets, failing to recover correct answers that were previously accessible to the base model. Interestingly, we also observe that while RLVR sometimes increases token-level entropy—resulting in greater uncertainty at each generation step—answer-level entropy declines, indicating that these seemingly more uncertain paths ultimately converge onto a smaller set of distinct answers. Taken together, these findings reveal potential limits of the current RLVR recipe in extending reasoning horizons. Breaking this invisible leash may require future algorithmic innovations such as explicit exploration mechanisms or hybrid strategies that seed probability mass into underrepresented solution regions.
1 Introduction
The rise of large reasoning models, such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024), marks a breakthrough in AI capabilities, particularly in solving complex logical tasks involving mathematics (Luo et al., 2025c; Zeng et al., 2025) and programming (Luo et al., 2025b; Liu & Zhang, 2025). The key ingredient behind this remarkable progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), where a pretrained base model—or one fine-tuned on long-form Chain-of-Thought (CoT) data—is optimized via reinforcement learning (RL) using simple, automatically computed rewards. Despite the empirical success, a fundamental question remains under active debate within the research community: does the current practice of RLVR expand a base model’s reasoning capabilities, or does it simply reinforce patterns the base model already knows, sometimes at the expense of exploring alternative correct solutions?
Recent empirical studies have revealed a puzzling pattern that hints at this limitation. While models trained with the common RLVR recipe consistently outperform their base models when evaluated with a single attempt (pass@1), the base models often perform better when given multiple attempts (pass@k for large k). It is even reported that RLVR-trained models benefit from seemingly random or spurious reward signals, raising questions about whether the observed improvements genuinely reflect enhanced reasoning (Shao et al., 2025). While pass@k may have limitations as a comprehensive measure of reasoning boundaries—as it primarily captures solution retrieval rather than novel reasoning capacity (Wen et al., 2025)—we adopt it here as a practical proxy metric following established practice in prior work (Chen et al., 2021; Shao et al., 2024; Liu et al., 2024; Chen et al., 2025b). This metric provides a useful lens for examining how RLVR affects solution accessibility, though future work should explore more nuanced measures of reasoning capability expansion.
Some researchers interpret this as evidence that the current RLVR recipe primarily performs conservative optimization within the base model’s existing capabilities (Yue et al., 2025a; Zhao et al., 2025; Shah et al., 2025; Ma et al., 2025; He et al., 2025). Others argue that this pattern only appears in specialized domains where base models were already well-trained, and that RLVR can substantially expand reasoning in other domains (Liu et al., 2025).
Figure 1: Empirical support dynamics of RLVR. Left: Conceptual illustration of empirical support under a threshold $\epsilon$. We define four regions based on whether a correct completion $y^{*}\in\mathcal{C}$ is assigned non-negligible probability mass by the base model $q$ and the RLVR model $\pi_{\theta}$: Empirical Support Preservation covers completions with $q(y^{*}|x)>\epsilon$ and $\pi_{\theta}(y^{*}|x)>\epsilon$; Empirical Support Shrinkage includes correct completions downweighted by RLVR below $\epsilon$; Empirical Support Expansion includes completions that RLVR newly upweights above $\epsilon$ despite negligible base model mass; and Out of Support refers to completions missed by both. Right: Pie charts showing the proportion of completions in each category across diverse reasoning tasks.
Seeking a definitive answer to this debate remains an open challenge. In the extreme case, it seems unlikely that the current RLVR recipe can unlock advanced reasoning capabilities for any model out of the box, such as GPT-2 (Radford et al., 2019). We are curious if there may exist inherent limitations in the current RLVR practice. This paper provides a systematic empirical investigation into the fundamental capabilities and potential limitations of the current RLVR practice. We introduce the concept of empirical support—the set of correct solutions that a model can realistically discover under finite sampling. Using this framework, we show that:
  1. 1.
    The current RLVR recipe primarily preserves rather than expands the base model’s solution coverage. Across diverse reasoning benchmarks, RLVR consistently loses access to more correct solutions than it gains, even while improving single-sample accuracy.
  2. 2.
    The precision-diversity trade-off is fundamental, not domain-specific. This pattern appears across mathematics, logical reasoning, factual QA, and code generation—suggesting it reflects inherent properties of current RLVR methods rather than domain-specific quirks.
  3. 3.
    Local uncertainty and global diversity can diverge. The current RLVR recipe sometimes increases token-level entropy (appearing more “uncertain” during generation) while simultaneously reducing answer-level entropy (converging to fewer final solutions).
These findings suggest that current RLVR methods may face an “invisible leash”—they remain fundamentally constrained by their initialization and cannot discover reasoning patterns that lie outside the base model’s effective reach. To break this invisible leash, RLVR may need augmenting with explicit exploration or hybrid strategies that seed probability mass into underrepresented regions of the solution space. We hope this work offers novel insights into the strengths and limitations of the current RLVR recipe, guiding future efforts in improving RLVR practice and building LLM systems that can unlock genuinely new reasoning capacity.
2 Preliminaries
2.1 Formalizing Solution Accessibility
Effective Support of Correct Completions.
Let $\mathcal{X}$ denote the space of natural language prompts, and $\mathcal{Y}$ denote the space of token sequences (e.g., reasoning traces or completions). For a fixed prompt $x\in\mathcal{X}$, $q(y\mid x)$ is the output distribution of the base model, and $R(x,y)\in\{0,1\}$ is a verifiable reward function indicating whether $y$ is a correct solution. Various RLVR algorithms, including PPO (Schulman et al., 2017), RLOO (Kool et al., 2019), GRPO (Guo et al., 2025), DAPO (Yu et al., 2025), or REINFORCE++ (Hu, 2025), learn a new distribution $\pi_{\theta}(y\mid x)$ to optimize different variants of the following regularized objective:
$\max_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x),x\sim\mathcal{D}}\left[R(x,y)-\beta^{-1}\log\frac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\right],$
where $\mathcal{D}$ is the distribution of prompts. An optional log ratio corresponds to a regularized policy update that penalizes divergence from the base model $q$ controlled by a hyperparameter $\beta>0$.
Definition 2.1 (Support of Correct Completions).
Let $\mathcal{C}=\{y\in\mathcal{Y}\mid R(x,y)=1\}$ denote the set of correct completions under the reward function $R$. Then the effective support on correct completions of a distribution $p(y\mid x)$ is defined as $\mathrm{supp}(p):=\left\{y\in\mathcal{C}\mid p(y\mid x)>0\right\}.$
Empirical Support Relaxation.
The effective support assumes that $q$ has exact zeros in its support, which, however, rarely holds in practice. Softmax layers yield strictly positive probabilities across all tokens, making the nominal support of $q$ span the entire space $\mathcal{Y}$. This factor, along with sampling noise or temperature scaling, contributes to what we refer to as empirical support diffusion: over time, the model may assign growing probability mass to completions that initially had negligible—but still nonzero—probability under the base model.
While $q(y\mid x)$ is technically positive for all $y$ due to the softmax, many completions lie so deep in the tail that they are effectively invisible to the training algorithm under finite sampling. To formalize this, we develop a relaxation and define the empirical support under $\epsilon$ as
$\mathrm{supp}_{\epsilon}(q):=\left\{y\in\mathcal{C}\mid q(y\mid x)>\epsilon\right\},$
where $\epsilon>0$, with $\epsilon\rightarrow 0$, denotes a minimal cutoff that separates completions with practically observable likelihood from those that are statistically negligible. Completions outside this threshold are unlikely to be sampled in typical on-policy RL settings with finite rollouts. The choice of $\epsilon$ is thus crucial for assessing which completions are empirically reachable. Intuitively, $\epsilon$ should correspond to the minimum probability required for a correct completion to appear within finite samples. We derive a principled estimate for this threshold based on sampling confidence bounds in Appx. C.4.
2.2 Characterizing How RLVR Changes Solution Access
With empirical support defined, we categorize what happens to correct solutions under RLVR:
Definition 2.2 (Empirical Support Dynamics).
For a given threshold $\epsilon>0$, We say RLVR achieves empirical support expansion under threshold $\epsilon$ if $\mathrm{supp}_{\epsilon}(\pi_{\theta})\setminus\mathrm{supp}_{\epsilon}(q)\neq\emptyset$, i.e. there exists at least one completion $y^{*}\in\mathcal{C}$ such that $q(y^{*}\mid x)\leq\epsilon\quad\text{but}\quad\pi_{\theta}(y^{*}\mid x)>\epsilon.$ That is, the RLVR-trained model assigns non-negligible probability mass to correct completions that were effectively negligible under the base model. We say RLVR exhibits empirical support shrinkage under threshold $\epsilon$ if $\mathrm{supp}_{\epsilon}(q)\setminus\mathrm{supp}_{\epsilon}(\pi_{\theta})\neq\emptyset$, i.e. there exists at least one completion $y^{*}\in\mathcal{C}$ such that $q(y^{*}\mid x)>\epsilon\quad\text{but}\quad\pi_{\theta}(y^{*}\mid x)\leq\epsilon.$ This formalizes the phenomenon where RLVR concentrates probability mass onto a narrower subset of outputs, effectively excluding correct solutions that were previously accessible under the base model.
Support Dynamics Metrics.
To quantify RLVR’s impact on solution accessibility, we introduce the following precision and recall-inspired metrics based on these four support categories.
Definition 2.3 (Support Dynamics Metrics).
Let $P$, $E$, $S$, and $O$ denote the number of correct completions in preservation, expansion, shrinkage, and out-of-support, respectively. Support Retention Rate (SRR) measures how well RLVR preserves the base model’s accessible correct solutions: $\text{SRR}=\frac{P}{P+S}$ Net Discovery Rate (NDR) measures the fraction of RLVR’s accessible solutions that represent genuine discoveries: $\text{NDR}=\frac{E}{P+E}$ Support Dynamic Score (SDS) provides a balanced measure combining retention and discovery: $\text{SDS}=\frac{2\cdot\text{SRR}\cdot\text{NDR}}{\text{SRR}+\text{NDR}}=\frac{2PE}{P^{2}+2PE+ES}$ Net Support Change Rate (NSCR) captures the net expansion or shrinkage of empirical support: $\text{NSCR}=\frac{E-S}{P+E+S}$
These metrics provide complementary perspectives on RLVR’s behavior:
  • SRR $\in[0,1]$: Higher values indicate better preservation of base model solutions. SRR =1=1 means no shrinkage occurred.
  • NDR $\in[0,1]$: Higher values indicate more genuine discovery. NDR =0=0 means no new solutions were found; NDR =1=1 means all accessible solutions are discoveries.
  • SDS $\in[0,1]$: Harmonic mean balancing retention and discovery. High SDS requires both good retention and meaningful expansion.
  • NSCR $\in[-1,1]$: Positive values indicate net expansion, negative values indicate net shrinkage.
These metrics enable us to distinguish between different RLVR behaviors: support-constrained optimization (high SRR, low NDR), genuine capability expansion (high SRR, high NDR), inefficient redistribution (low SRR, low NDR), and aggressive exploration (low SRR, high NDR). We also provide theoretical foundations to understand RLVR’s support-bounded behavior in Appx. C.
3 Evidence of Hidden-Support Dynamics
3.1 Experimental Setup
We adopt ProRL-1.5B-v1 (Liu et al., 2025) as our main RLVR method due to its robust long-horizon training framework. Starting from DeepSeek-R1-Distill-Qwen-1.5B as the base model, ProRL’s Nemotron-Research-Reasoning series leverages GRPO enhanced with decoupled clipping, dynamic sampling, KL divergence regularization, and periodic reference resets to sustain exploration and prevent entropy collapse during extended RL training. In addition, we evaluate other RLVR variants at multiple scales (7B–14B parameters), including Skywork (Wei et al., 2023), AceReason-Nemotron (Chen et al., 2025a), and Phi4-Reason (Abdin et al., 2025), alongside a visual LLM (Kangheng-OVR-7B (Wei et al., 2025)).
Performance is measured across two categories. (1) Math reasoning tasks: MATH500 (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), AIME 2024, AIME 2025, and AMC 2023. (2) Non-math reasoning tasks: SimpleQA (Wei et al., 2024) (factuality), LiveBench (White et al., 2025) (logic, coding, and language comprehension), SciBench (Wang et al., 2023) (multi-domain scientific problem-solving), and Reasoning Gym (Stojanovski et al., 2025) (cognition, geometry, graph theory, games). In Reasoning Gym, we especially focus on tasks that ProRL explicitly highlighted as challenging for the base model. For SimpleQA, we employ GPT-4.1 (Achiam et al., 2023) as the judge. Sampling budgets are $k\in\{4096,8192\}$ for math, $k\in\{1024,2048,4096,8192,16384\}$ for Reasoning Gym, and $k\in\{1024,2048\}$ for other non-math datasets, ensuring that any unreachable solution $y^{*}\in\mathcal{C}$ remains below the empirical support threshold of the base model. More implementation details appear in Appx. B.
Table 1: Aggregate support dynamics across diverse models and domains. Each completion is categorized by correctness and support status: Preservation indicates both base and RLVR find the solution; Shrinkage indicates the base model found it but RLVR did not; Expansion indicates only RLVR found it; and Out of Support denotes solutions found by neither. Higher SRR, NDR, and SDS reflect stronger preservation, genuine discovery, and balanced optimization, respectively. NSCR values closer to zero indicate more balanced support change. Kangheng-OVR-7B is included as a vision-language model (VLM). Full detailed statistics for each model are provided in Appx. A.
Model Domain SRR NDR SDS NSCR P E S O
ProRL-1.5B-v1 Math 0.96 0.00 0.01 -0.04 1355 5 56 131
Non-Math 0.91 0.03 0.06 -0.06 1045 31 107 674
Overall 0.94 0.02 0.03 -0.05 2400 36 163 805
ProRL-1.5B-v2 Math 0.96 0.01 0.01 -0.04 1349 9 62 127
Non-Math 0.90 0.04 0.07 -0.06 1039 39 113 666
Overall 0.93 0.02 0.04 -0.05 2388 48 175 793
Nemotron-1-7B Math 0.99 0.00 0.01 -0.00 1431 5 9 102
Non-Math 0.97 0.02 0.04 -0.02 1284 23 47 503
Overall 0.98 0.01 0.02 -0.01 2715 28 56 605
Skywork-OR1-7B Math 0.98 0.00 0.00 -0.02 1406 2 34 105
Non-Math 0.96 0.02 0.04 -0.02 1279 24 52 502
Overall 0.97 0.01 0.02 -0.02 2685 26 86 607
Nemotron-1-14B Math 0.99 0.00 0.01 -0.01 1425 5 15 102
Non-Math 0.99 0.00 0.01 -0.01 993 3 8 399
Overall 0.99 0.00 0.01 -0.01 2418 8 23 501
Phi4-Reason-Plus-14B Math 0.99 0.01 0.01 -0.00 1407 8 12 120
Non-Math 0.99 0.01 0.01 -0.00 1067 8 11 317
Overall 0.99 0.01 0.01 -0.00 2474 16 23 437
Kangheng-OVR-7B (VLM) Math 1.00 0.00 0.01 -0.00 781 3 4 516
Figure 2: Typical empirical support preservation in Reasoning Gym tasks, like Graph Coloring, Palindrome Generation, and Advanced Geometry.
Figure 3: Instances of empirical support expansion, as seen in Boxnet, Dice, and Arc 1D tasks.
Figure 4: Examples of empirical support shrinkage on Reasoning Gym tasks such as Leg Counting, Family Relationships, and Power Function.
3.2 Results: Predominant Preservation with Limited Expansion
Support preservation dominates across all domains.
Table 1 shows that across diverse model scales and families, RLVR predominantly acts as a support-constrained optimization mechanism. All models achieve very high support retention (overall $\text{SRR}\approx 0.93$$0.99$) while genuine discovery remains rare ($\text{NDR}\leq 0.04$). For example, Nemotron-7B and Nemotron-14B retain nearly all base-model solutions ($\text{SRR}\geq 0.98$) with negligible expansion. Even smaller-scale ProRL-1.5B achieves $\text{SRR}=0.93$ with only modest gains ($\text{NDR}=0.02$). These patterns persist across math ($\text{SRR}=0.96$$0.99$, $\text{NDR}\approx 0.00$$0.01$) and non-math domains ($\text{SRR}=0.90$$0.99$, $\text{NDR}\leq 0.04$). Preservation-dominated behavior is especially clear in Reasoning Gym tasks such as graph_color and palindrome, where ProRL accelerates convergence toward near-perfect pass@k with large budgets (Fig. 2). Support counts confirm this: most correct completions remain shared between RLVR and base models.
Selective but limited empirical support expansion.
Despite the strong conservation, RLVR occasionally recovers solutions negligible to the base model. Expansion is consistently small: ProRL-1.5B discovers 48 new completions across 11 benchmarks, while larger models (e.g., Phi4-14B, Nemotron-14B) add fewer than 10. Non-math datasets exhibit the highest relative discovery ($\text{NDR}\leq 0.04$), whereas math datasets are virtually stagnant ($\text{NDR}\leq 0.01$). Some Reasoning Gym tasks, such as graph_color_vertex20 and arc_1d, show genuine expansion (Fig. 3), but remain isolated exceptions rather than the dominant trend. These suggest that while RLVR can occasionally redistribute mass into underexplored solution modes, such expansion remains the exception rather than the rule, challenging assumptions about RLVR’s capacity for genuine reasoning horizon extension.
Empirical support shrinkage outweighs expansion.
Across all models and domains, shrinkage consistently exceeds expansion. ProRL-1.5B loses 175 completions while gaining only 48 (ratio $\approx 3.6{:}1$), while Nemotron-7B and Skywork-OR1-7B display similar patterns (ratios $\approx 2{:}1$3:13{:}1). Even large models (Nemotron-14B, Phi4-14B) show net shrinkage despite near-perfect preservation. Overall NSCR values remain slightly negative (0.01-0.01 to 0.06-0.06), showing that RLVR systematically narrows the accessible solution set. This explains paradoxical outcomes: while RLVR models outperform bases at low $k$, base models dominate at high $k$ due to broader solution coverage (e.g., AIME2024 base $\texttt{pass@8192}=93.3\%$ vs. ProRL-1.5B’s $83.3\%$). Reasoning Gym tasks like leg_counting, family_relationships, and power_function illustrate this vividly (Fig. 4).
Support dynamic score confirms imbalance. SDS values remain consistently low across all scales ($\leq 0.07$), reflecting poor balance between preservation and discovery. The highest observed SDS is LiveBench-L at $0.288$, but even this corresponds to 2 expansions against 4 shrinkages. Math benchmarks are particularly imbalanced (SDS $\approx 0.00$$0.01$), while non-math domains fare only marginally better. Thus, RLVR’s improvements arise primarily from mass concentration, not meaningful solution expansion.
Perplexity analysis on support constraints. Tab. 2 reports perplexity scores under two complementary settings. When evaluated against external reasoning traces from DeepSeek-R1 and Claude Sonnet 4 with extended thinking—likely outside the base model’s support—RLVR shows markedly higher perplexity (e.g., AIME2024 rising from $8.76$ to $14.91$), confirming its inability to assign mass to fundamentally novel solution modes. Differences in language style and reasoning format across external references (e.g., Claude vs DeepSeek) also contribute to perplexity gaps, beyond purely structural support constraints. Breakdowns by correctness patterns highlight the precision–coverage trade-off: in shrinkage cases, ProRL’s perplexity rises when it fails to recover solutions still accessible to the base, reflecting entropy-driven concentration. Meanwhile, modest perplexity gaps in rare expansion cases indicate that these new completions were drawn from the base’s long-tail low-density regions—amplified but not truly beyond its support.
Overall takeaway: RLVR as precision enhancer, not capability expander. Across model scales (1.5B–14B) and domains (math, non-math, multimodal), RLVR consistently behaves as a support-bounded optimizer. With SRR near one but NDR near zero, and uniformly negative NSCR, RLVR enhances precision by concentrating mass on known high-reward solutions but rarely discovers new reasoning paths. This aligns with the Temporal Forgetting effect (Li et al., 2025). Breaking RLVR’s invisible leash may thus require explicit exploration or hybrid strategies that deliberately seed probability mass into underrepresented solution regions.
Table 2: Perplexity of reasoning tokens from base and RLVR across math benchmarks, segmented by correctness patterns and reference types. Top: on problems correctly solved by the base model but not RLVR, perplexity is measured against the base model’s reasoning traces. Middle: on problems correctly solved by RLVR but not the base, perplexity is measured against RLVR’s traces. Bottom: on problems unsolved by both, perplexity is computed against external references (DeepSeek-R1 and Claude Sonnet 4), reflecting each model’s compatibility with broader solution modes.
Category Correctness Reference Target AIME 2024 AIME 2025 Olympiad
Shrinkage ✓ Base, ✗ ProRL Base Base 1.36 1.47 1.30
ProRL 1.60 1.84 1.50
Expansion ✗ Base, ✓ ProRL ProRL Base - - 1.52
ProRL - - 1.32
– – ✗ Base, ✗ ProRL DeepSeek-R1 Base 1.82 1.75 1.62
ProRL 2.20 2.15 1.94
Claude Sonnet 4 Base 8.76 6.05 5.98
ProRL 14.91 9.76 9.55
4 Entropy Reduction and the pass@k Trade-off
4.1 Experimental Setup
To study how RLVR reshapes the sampling distribution, we examine the base model and RLVR with a medium sampling budget $k=32$ on the math reasoning benchmarks. We quantify changes in the output distribution using two entropy metrics:
  • Token-Level Entropy: Let $\mathcal{V}$ denote the vocabulary and $y^{(i)}=(y^{(i)}_{1},y^{(i)}_{2},\ldots,y^{(i)}_{T^{(i)}})$ denote the $i$-th generated sequence of length $T^{(i)}$ for $1\leq i\leq N$. At each timestep $t$, the model outputs a probability distribution $p_{t}^{(i)}(v)$ over vocabulary tokens $v\in\mathcal{V}$. The entropy of this distribution is given by: $H_{t}^{(i)}=-\sum_{v\in\mathcal{V}}p_{t}^{(i)}(v)\log p_{t}^{(i)}(v)$. The average token-level entropy over all $N$ sequences and their timesteps is computed as: $\text{TokenEntropy}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{T^{(i)}}\sum_{t=1}^{T^{(i)}}H_{t}^{(i)}\right)$, capturing the local uncertainty at each generation step.
  • Answer-level Entropy: Let $\{o^{(1)},\ldots,o^{(N)}\}$ denote the answers extracted from each generated sequence $y^{(i)}$ (using NA for incomplete outputs), and let $\{o_{1}^{\ast},\ldots,o_{M}^{\ast}\}$ be the $M$ unique answers. Let $f_{j}$ be the frequency of answer $o_{j}^{\ast}$, with empirical probability $p_{j}=\frac{f_{j}}{N}$. Then: $\text{AnswerEntropy}=-\sum_{j=1}^{M}p_{j}\log p_{j}$. This captures global diversity over output completions, with lower values indicating increased mode collapse.
4.2 Results: Precision Gains, Entropy Dynamics, and Trade-offs
Table 3: Summary of avg@32 accuracy, response length, and entropy metrics across math reasoning benchmarks (row colors: $x$ base models, $x$ RLVR models). RLVR consistently improves accuracy and alters distributional properties. While answer-level entropy consistently decreases, token-level entropy shows more varied behavior across models.
Metric Model AIME 2024 AMC 2023 MATH 500 Minerva Olympiad Avg.
DeepSeek-1.5B 31.15 72.81 85.01 32.18 51.55 54.54
ProRL-1.5B 45.62 85.70 92.01 39.27 64.56 65.43
DeepSeek-7B 53.23 89.30 93.95 43.07 66.67 69.24
avg@32 AceReason-7B 65.83 95.08 95.81 45.35 73.92 75.20
Acc. (%) Skywork-OR1-7B 67.40 93.59 95.73 43.81 73.05 74.71
DeepSeek-14B 67.81 95.39 95.28 46.43 72.06 75.39
AceReason-14B 77.29 98.67 97.01 47.20 77.74 79.58
Qwen2.5-32B 18.12 55.23 75.84 24.55 41.40 43.03
DAPO-32B 51.25 92.81 80.75 32.50 49.15 61.29
DeepSeek-1.5B 16363 9979 5700 8194 11873 10422
ProRL-1.5B 7786 6294 5070 6569 6678 6479
DeepSeek-7B 13613 6402 4125 5595 8988 7745
Response AceReason-7B 10740 5961 4313 6261 7703 6995
Length Skywork-OR1-7B 15628 $82$ 5735 8742 12094 10096
DeepSeek-14B 11295 5735 3781 4919 8042 6755
AceReason-14B 13871 7239 4622 7720 10033 8697
Qwen2.5-32B 1247 874 585 3544 881 1426
DAPO-32B 6908 3157 3386 5665 5827 4989
DeepSeek-1.5B 0.45 0.40 0.42 0.49 0.44 0.44
[…]
DeepSeek-7B 0.38 0.34 0.35 0.39 0.38 0.37
[…]
DeepSeek-14B 0.33 0.30 0.32 0.35 0.33 0.33
[…]
Qwen2.5-32B 0.17 0.16 0.15 0.28 0.15 0.18
[…]
DeepSeek-1.5B 2.15 0.91 0.46 1.65 1.33 1.30
ProRL-1.5B 1.24 0.35 0.18 0.90 0.63 0.66
DeepSeek-7B 1.47 0.36 0.18 0.96 0.80 0.75
Answer-Level AceReason-7B 0.96 0.14 0.11 0.77 0.53 0.50
Entropy Skywork-OR1-7B 0.97 0.20 0.12 0.80 0.58 0.54
DeepSeek-14B 1.01 0.14 0.13 0.83 0.59 0.54
AceReason-14B 0.66 0.06 0.07 0.67 0.44 0.38
Qwen2.5-32B 2.37 1.32 0.68 2.27 1.41 1.61
DAPO-32B 1.12 0.09 0.26 0.96 0.63 0.61
Consistent gains in precision, but sharper global distributions.
Tab. 3 shows that RLVR consistently improves avg@32 across all benchmarks, raising average performance from 54.5% to 65.4% for ProRL and from 43.0% to 61.3% for DAPO (Yu et al., 2025). However, this increased precision comes at a cost: RLVR systematically reduces answer-level entropy, indicating a collapse onto fewer distinct solutions and empirically validating our theoretical prediction that reward optimization sharpens output distributions around known modes, thereby reducing effective support coverage. Notably, intrinsically harder tasks, such as AIME or Minerva, still exhibit higher absolute answer-level entropy for both the base and RLVR models, suggesting that challenging problems inherently foster broader solution spaces that require exploration over more diverse completions.
Decoupled local uncertainty and global diversity.
While answer-level entropy consistently declines, token-level entropy exhibits more varied behavior. In models like ProRL and DAPO, it increases, suggesting greater local uncertainty during generation, possibly due to longer or more elaborated reasoning chains that introduce additional decision points or “forking” tokens (Wang et al., 2025). However, this pattern is far from universal: other RLVR models like AceReason and Skywork display similar or even lower token-level entropy relative to their base counterparts, and prior work has documented sharp entropy collapse in early training phases (Cui et al., 2025).
More importantly, increased token-level entropy does not imply greater exploration of the output space. Despite appearing more stochastic at the step level, RLVR models frequently converge onto a smaller set of final answers—reflected in lower answer-level entropy. Notably, even between two models built on the same base (DeepSeek-7B), Skywork-OR1-7B shows lower token-level entropy than AceReason-7B, yet exhibits higher answer-level entropy. This contrast highlights that local uncertainty does not reliably predict the diversity of final solutions, revealing a critical decoupling between local uncertainty and global diversity. We refer to this phenomenon as local stochasticity without global exploration: the model exhibits variability in generation but ultimately collapses to a narrow set of solutions. Thus, token-level entropy should not be conflated with genuine exploratory behavior, and interpreting entropy dynamics in RLVR requires distinguishing between stepwise uncertainty and overall support expansion.
Implications.
Our empirical analysis reveals a trade-off in RLVR: it improves precision by amplifying high-reward outputs, but simultaneously narrows the diversity of global solutions. This limitation is especially consequential in domains that admit multiple valid answers or benefit from creative reasoning, underscoring the need for explicit exploration mechanisms or diversity-promoting strategies to complement standard RLVR. Moreover, the observed divergence between token-level and answer-level entropy highlights the need for a more nuanced interpretation of stochasticity in reward-optimized models—showing that precision gains often come at the expense of global diversity, and that maintaining controlled variability is critical for sustaining effective exploration.
5 Conclusion
We reveal that the current RLVR improves precision by sharpening distributions around known high-reward trajectories, yet largely preserves the base model’s support. Importantly, we found that this sharpening does not merely prune incorrect outputs—it can also concentrate probability mass on a narrower subset of correct solutions, occasionally excluding valid alternatives that the more diverse base model could still recover. Meanwhile, the divergence between token-level uncertainty and answer-level diversity indicates that stepwise stochasticity alone is insufficient for global exploration, motivating future work to bridge this gap. We suggest that to expand reasoning capabilities beyond the base model’s scope, RLVR must be coupled with explicit exploration strategies or off-policy mechanisms that seed probability mass into underrepresented regions of the solution space.
Ethics Statement
This research was conducted in accordance with established academic ethical standards using publicly available models and appropriate computational resources. All models were accessed through proper channels with necessary permissions, and outputs were analyzed solely for research purposes to understand fundamental properties of reinforcement learning techniques. We acknowledge the significant computational resources required for LLM evaluation and made efforts to optimize experimental design while maintaining scientific rigor. This work investigates limitations of current RL approaches in AI systems, and we believe transparent reporting of these constraints is essential for guiding effective future research directions, helping practitioners set appropriate expectations, and contributing to broader scientific understanding of AI capabilities. While our analysis focuses on specific model families and current benchmarks, which may limit generalizability, we encourage continued research into these limitations and the development of methods that can genuinely expand model reasoning capabilities while maintaining the beneficial precision improvements that RLVR provides.
Appendix A Detailed Statistics for Support Dynamics
We provide full per-model statistics that underlie the aggregate values in Table 1. For each model and domain (Math, Non-Math, and Overall), we report the raw counts of correct completions across the four empirical support categories: Preservation (P), Expansion (E), Shrinkage (S), and Out-of-Support (O). From these counts, we compute the derived metrics: Support Retention Rate (SRR), Net Discovery Rate (NDR), Support Dynamic Score (SDS), and Net Support Change Rate (NSCR).
These expanded tables enable a fine-grained comparison of how different RLVR variants and model scales redistribute probability mass across correct solutions. In particular, they clarify whether improvements in single-sample accuracy stem from strong preservation of the base model’s support, from genuine discovery of new solutions, or from trade-offs between expansion and shrinkage.
We include results for all evaluated models: ProRL-1.5B-V2, Nemotron-1-7B, Skywork-OR1-7B, AceReason-Nemotron-1-14B, Phi4-Reason-Plus-14B, and the visual reasoning model Kangheng-OVR-7B (VLM). These tables serve as the ground truth for the aggregate summaries in the main text and substantiate the claims about predominant preservation, limited expansion, and consistent shrinkage observed across both math and non-math domains.
Table 4: Support dynamics metrics and pass@k performance of ProRL-1.5B-v1, compared with its base model, DeepSeek-R1-Distill-Qwen-1.5B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@8192)
AIME2024 93.3% 83.3% 0.893 0.000 0.000 -0.107 25 0 3 2
AIME2025 80.0% 73.3% 0.833 0.091 0.164 -0.077 20 2 4 4
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 99.6% 99.4% 0.998 0.000 0.000 -0.002 497 0 1 2
Minerva 71.7% 63.6% 0.887 0.000 0.000 -0.113 173 0 22 77
Olympiad 92.7% 89.3% 0.958 0.005 0.010 -0.037 600 3 26 46
Non-Math Reasoning Benchmarks (pass@2048)
SimpleQA 23.3% 18.0% 0.743 0.038 0.073 -0.221 75 3 26 329
LiveBench-R 100.0% 94.0% 0.940 0.000 0.000 -0.060 94 0 6 0
LiveBench-C 62.5% 56.2% 0.838 0.069 0.128 -0.094 67 5 13 43
LiveBench-L 26.0% 24.0% 0.769 0.167 0.274 -0.067 10 2 3 35
SciBench 94.1% 90.5% 0.946 0.016 0.031 -0.038 616 10 35 31
LiveCodeBench v5 46.4% 43.0% 0.860 0.072 0.133 -0.069 129 10 21 163
LiveCodeBench v6 43.5% 42.0% 0.947 0.018 0.036 -0.034 54 1 3 73
Aggregate Statistics
Math Benchmarks 0.960 0.0037 0.0073 -0.036 1355 5 56 131
Non-Math Benchmarks 0.907 0.0288 0.0558 -0.064 1045 31 107 674
Overall 0.936 0.0148 0.0291 -0.049 2400 36 163 805
Table 5: Support dynamics metrics and pass@k performance of ProRL-1.5B-v2, compared with its base model, DeepSeek-R1-Distill-Qwen-1.5B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@8192)
AIME2024 93.3% 90.0% 0.964 0.000 0.000 -0.036 27 0 1 2
AIME2025 80.0% 73.3% 0.875 0.045 0.086 -0.080 21 1 3 5
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 99.6% 99.6% 1.000 0.000 0.000 0.000 498 0 0 2
Minerva 71.7% 63.2% 0.856 0.029 0.056 -0.115 167 5 28 72
Olympiad 92.7% 88.7% 0.952 0.005 0.010 -0.043 596 3 30 46
Non-Math Reasoning Benchmarks (pass@2048)
SimpleQA 23.3% 22.2% 0.822 0.135 0.233 -0.044 83 13 18 319
LiveBench-R 100.0% 92.0% 0.920 0.000 0.000 -0.080 92 0 8 0
LiveBench-C 62.5% 60.2% 0.887 0.078 0.143 -0.035 71 6 9 42
LiveBench-L 26.0% 22.0% 0.692 0.182 0.288 -0.133 9 2 4 35
SciBench 94.1% 88.9% 0.937 0.008 0.016 -0.055 610 5 41 36
LiveCodeBench v5 46.4% 41.2% 0.820 0.075 0.138 -0.106 123 10 27 163
LiveCodeBench v6 43.5% 41.2% 0.895 0.056 0.105 -0.050 51 3 6 71
Aggregate Statistics
Math Benchmarks 0.956 0.007 0.013 -0.037 1349 9 62 127
Non-Math Benchmarks 0.902 0.036 0.070 -0.062 1039 39 113 666
Overall 0.932 0.020 0.039 -0.049 2388 48 175 793
Table 6: Support dynamics metrics and pass@k performance of AceReason-Nemotron-1-7B, compared with its base model, DeepSeek-7B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@8192)
AIME2024 93.3% 93.3% 1.000 0.000 0.000 0.000 28 0 0 2
AIME2025 100.0% 100.0% 1.000 0.000 0.000 0.000 30 0 0 0
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 99.8% 99.8% 1.000 0.000 0.000 0.000 499 0 0 1
Minerva 71.7% 71.0% 0.985 0.005 0.010 -0.010 192 1 3 76
Olympiad 96.0% 95.7% 0.991 0.006 0.012 -0.003 642 4 6 23
Non-Math Reasoning Benchmarks (pass@2048)
SimpleQA 38.6% 35.6% 0.862 0.065 0.121 -0.073 144 10 23 256
LiveBench-R 100.0% 99.0% 0.990 0.000 0.000 -0.010 99 0 1 0
LiveBench-C 85.9% 85.9% 0.991 0.009 0.018 0.000 109 1 1 17
LiveBench-L 24.0% 24.0% 0.833 0.167 0.278 0.000 10 2 2 36
SciBench 94.7% 93.5% 0.982 0.006 0.012 -0.012 643 4 12 33
LiveCodeBench v5 62.8% 62.5% 0.970 0.025 0.048 -0.005 197 5 6 115
LiveCodeBench v6 64.1% 63.4% 0.976 0.012 0.024 -0.012 82 1 2 46
Aggregate Statistics
Math Benchmarks 0.994 0.003 0.007 -0.003 1431 5 9 102
Non-Math Benchmarks 0.965 0.018 0.035 -0.018 1284 23 47 503
Overall 0.981 0.010 0.020 -0.010 2715 28 56 605
Table 7: Support dynamics metrics and pass@k performance of Skywork-OR1-7B, compared with its base model, DeepSeek-R1-Distill-Qwen-7B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@8192)
AIME2024 93.3% 93.3% 1.000 0.000 0.000 0.000 28 0 0 2
AIME2025 100.0% 100.0% 1.000 0.000 0.000 0.000 30 0 0 0
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 99.8% 99.8% 1.000 0.000 0.000 0.000 499 0 0 1
Minerva 71.7% 71.3% 0.985 0.010 0.020 -0.005 192 2 3 75
Olympiad 96.0% 91.4% 0.952 0.000 0.000 -0.048 617 0 31 27
Non-Math Reasoning Benchmarks (pass@2048)
SimpleQA 38.6% 37.0% 0.880 0.081 0.149 -0.039 147 13 20 253
LiveBench-R 100.0% 98.0% 0.980 0.000 0.000 -0.020 98 0 2 0
LiveBench-C 85.9% 85.2% 0.991 0.000 0.000 -0.009 109 0 1 18
LiveBench-L 24.0% 22.0% 0.917 0.000 0.000 -0.083 11 0 1 38
SciBench 94.7% 92.8% 0.974 0.006 0.012 -0.020 638 4 17 33
LiveCodeBench v5 62.8% 62.2% 0.966 0.025 0.049 -0.010 196 5 7 115
LiveCodeBench v6 64.1% 62.6% 0.952 0.024 0.048 -0.023 80 2 4 45
Aggregate Statistics
Math Benchmarks 0.976 0.001 0.003 -0.023 1406 2 34 105
Non-Math Benchmarks 0.961 0.018 0.036 -0.021 1279 24 52 502
Overall 0.969 0.010 0.020 -0.022 2685 26 86 607
Table 8: Support dynamics metrics and pass@k performance of Nemotron-1-14B, compared with its base model, DeepSeek-R1-Distill-Qwen-14B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@4096)
AIME2024 96.7% 93.3% 0.966 0.000 0.000 -0.034 28 0 1 1
AIME2025 100.0% 96.7% 0.967 0.000 0.000 -0.033 29 0 1 0
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 99.8% 99.8% 1.000 0.000 0.000 0.000 499 0 0 1
Minerva 71.7% 69.5% 0.959 0.011 0.021 -0.030 187 2 8 75
Olympiad 95.9% 95.6% 0.992 0.005 0.009 -0.003 642 3 5 25
Non-Math Reasoning Benchmarks (pass@1024)
SimpleQA 27.0% 26.8% 0.983 0.009 0.017 -0.008 115 1 2 315
LiveBench-R 99.0% 99.0% 1.000 0.000 0.000 0.000 99 0 0 1
LiveBench-C 92.2% 92.2% 1.000 0.000 0.000 0.000 118 0 0 10
LiveBench-L 46.0% 44.0% 0.957 0.000 0.000 -0.043 22 0 1 27
SciBench 93.1% 92.6% 0.992 0.003 0.006 -0.005 639 2 5 46
Aggregate Statistics
Math Benchmarks 0.990 0.0035 0.0070 -0.0069 1425 5 15 102
Non-Math Benchmarks 0.992 0.0030 0.0060 -0.0050 993 3 8 399
Overall 0.991 0.0033 0.0066 -0.0061 2418 8 23 501
Table 9: Support dynamics metrics and pass@k performance of Phi4-Reason-Plus-14B, compared with its base model – Phi4-Reason-14B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@4096)
AIME2024 96.7% 96.7% 1.000 0.000 0.000 0.000 29 0 0 1
AIME2025 100.0% 100.0% 1.000 0.000 0.000 0.000 30 0 0 0
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 100.0% 99.8% 0.998 0.000 0.000 -0.002 499 0 1 0
Minerva 66.2% 65.4% 0.972 0.017 0.033 -0.011 175 3 5 89
Olympiad 94.8% 94.7% 0.991 0.008 0.016 -0.002 634 5 6 30
Non-Math Reasoning Benchmarks (pass@1024)
SimpleQA 37.9% 37.4% 0.970 0.019 0.036 -0.012 159 3 5 266
LiveBench-R 100.0% 100.0% 1.000 0.000 0.000 0.000 100 0 0 0
LiveBench-C 97.7% 96.9% 0.992 0.000 0.000 -0.008 124 0 1 3
LiveBench-L 74.0% 74.0% 1.000 0.000 0.000 0.000 37 0 0 13
SciBench 94.2% 94.2% 0.992 0.008 0.015 0.000 647 5 5 35
Aggregate Statistics
Math Benchmarks 0.992 0.0057 0.0112 -0.0028 1407 8 12 120
Non-Math Benchmarks 0.990 0.0074 0.0148 -0.0028 1067 8 11 317
Overall 0.991 0.0064 0.0128 -0.0028 2474 16 23 437
Table 10: Support dynamics metrics and pass@k performance of vision-language model OVR-7B-RL, compared with its base model, OVR-7B-ColdStart, across visual math reasoning benchmarks.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@8192)
MathVista 49.1% 49.1% 0.998 0.002 0.004 0.000 490 1 1 508
MathVision 96.7% 96.4% 0.990 0.007 0.014 -0.003 291 2 3 8
Aggregate Statistics
Math Benchmarks 0.995 0.0038 0.0076 -0.0013 781 3 4 516
Appendix B Experimental Details
We provide comprehensive details of the experimental setup, including dataset descriptions and evaluation methodologies. A key aspect of our evaluation approach is the answer processing enhancement framework for Reasoning Gym, which addresses format compatibility challenges between base and ProRL models to ensure fair evaluation.
B.1 Evaluation Settings
We employed vLLM (Kwon et al., 2023) as the inference backend. For all models, we utilized a sampling temperature of $0.6$, a top_ptop\_p value of $0.95$, and a maximum response length of $32768$.
B.2 Datasets
Math benchmarks.
We utilized the complete datasets from MATH500 (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), AIME 2024, AIME 2025, and AMC 2023 for evaluating LLMs. For vision-language models, we evaluated on the testmini sets of MathVision (Wang et al., 2024) and MathVista (Lu et al., 2023).
Non-math benchmarks.
For SimpleQA (Wei et al., 2024), we uniformly sampled 10% of the original dataset (433 samples) to enable efficient large-scale evaluation under high-pass conditions. For LiveBench (White et al., 2025), we used the 2024-11-25 version available on HuggingFace. To ensure unambiguous evaluation, we focused exclusively on tasks with binary correct/incorrect judgments and excluded tasks involving intermediate floating-point judgments, as these lack clear correctness criteria. Based on this selection criterion, we evaluated the following subsets: web_of_lies_v2web\_of\_lies\_v2 and spatialspatial subsets for Reasoning tasks (LiveBench-R), the typostypos subset for Language tasks (LiveBench-L), and all available data for Coding tasks (LiveBench-C). For SciBench (Wang et al., 2023), we evaluated on the complete dataset. For LiveCodeBench (Jain et al., 2024), we evaluated the dataset on both v5 and v6 versions. Due to computational efficiency considerations, we conducted LiveCodeBench evaluation exclusively on 1.5B and 7B models, excluding the 14B variants from this particular benchmark.
Reasoning Gym.
For Reasoning Gym (Stojanovski et al., 2025), we employ the easy set from the version updated after commit 17a8431 in its repository as our default task configuration. This choice ensures consistency with the default task configuration used in prior evaluations, maintaining comparable experimental conditions. Additionally, we utilize the hard set as our challenging evaluation benchmark for further evaluations.
B.3 Answer Processing Enhancement in Reasoning Gym
We identified significant evaluation challenges when testing the base model on Reasoning Gym. The ProRL model, having been trained on Reasoning Gym data, predominantly produces responses that conform to the expected format, leading to much higher accuracy scores. In contrast, the base model struggled with format adherence due to insufficiently detailed prompts, and its limited 1.5B parameter capacity made it particularly susceptible to evaluation inconsistencies. To address these issues, we enhanced both the answer extraction protocol and prompt design to ensure fair and objective accuracy assessments across both models. This causes the differences of ProRL’s reported performance and our evaluation results in Reasoning Gym.
B.3.1 General Answer Extraction Protocol
First, we enhanced the answer extraction protocol with a hierarchical, priority-based extraction mechanism that processes responses through multiple fallback levels. Each level attempts to capture the model’s intended answer, and successful extraction at any level bypasses subsequent processing steps.
[…]
As a final fallback, the system extracts content following </think> tags when all previous methods fail and the response contains these markers. This safety mechanism captures base model responses that ignore formatting requirements in lengthy tasks.
B.3.2 Task-Specific Processing Modifications
Our core answer processing pipeline applies to both models, with additional processing steps designed primarily to address format compatibility issues commonly encountered with base model responses. Specifically, the processing logic for each task is enhanced as follows:
dicedice
The ground truth for dice tasks uses $a/b$ fraction format. Base models frequently express fractions in format, requiring format standardization for accurate evaluation. For base models only, we convert ground truth fractions from $a/b$ format to format $\frac${a}{b} to ensure both model answers and ground truth use consistent formatting. ProRL dice processing maintains $a/b$ formatting for both model answers and ground truth, leveraging the dice samples present in its training data.
prime_factorizationprime\_factorization
The ground truth format requires answers to be combinations of numbers and multiplication symbol (i.e., $\times$) only. We implement three key modifications to ensure compatibility with this requirement. First, we standardize multiplication symbols by replacing $\times$ with $\times$ to meet the evaluation requirements, as base models frequently use multiplication symbols instead of standard multiplication signs. Second, we expand exponentiation by converting formats like a^b into repeated multiplication ($a\times a\times\ldots\times a$ for $b$ iterations), preventing errors when base models consolidate repeated factors into exponential notation. Third, we process response equations by retaining only right-side content when answers contain equals signs, transforming responses like “561 = 3 $\times$ 11 $\times$ 17” to “3 $\times$ 11 $\times$ 17” to eliminate question restatement that base models commonly include.
palindrome_generationpalindrome\_generation
The ground truth format expects palindromic character strings (sequences that read the same forwards and backwards). We remove excess whitespace to address frequent spacing issues in base model responses. This transformation converts spaced responses like “k h g a g h k” to “khgaghk”, preventing string reversibility judgment failures that occur when spaces interfere with palindrome verification.
advanced_geometryadvanced\_geometry
[…]
power_functionpower\_function
The ground truth format uses e-notation scientific notation. We convert mixed and arithmetic symbol scientific notation to ensure format consistency. The system transforms patterns like $2.36\times 10^{-16}$” or $1.5\times 10^{5}$” to e-notation format (“-2.36e-16”, “1.5e5”), preventing numerically correct but format-incompatible evaluation errors when base models use mixed and arithmetic symbols for scientific notation.
arc_1darc\_1d
[…]
boxnetboxnet
The ground truth format requires dictionary list formatting [{key: value}, ...]. We implement comprehensive JSON format cleaning to meet these evaluation requirements. Our processing pipeline includes several steps: rejecting pure numerical responses to prevent non-JSON format interference; removing JSON markdown wrappers that eliminate ‘‘‘json {content} ‘‘‘ markers; converting single dictionaries to single-element dictionary lists (dict $\rightarrow$ [dict]); and filtering illegal elements by removing non-dictionary components from JSON lists. Additionally, we clean nested structure values within individual dictionary entries. For nested lists, we extract the first element as the value ([{key1: [value1, value2, ...]}, ...] $\rightarrow$ [{key1: value1}, ...]). For nested dictionaries, we select matching key values when available ([{key1: {key1: value1, key2: value2, ...}}, ...] $\rightarrow$ [{key1: value1}, ...]) or default to the first element value when keys don’t match ([{key1: {key2: value2, key3: value3}}, ...] $\rightarrow$ [{key1: value2}, ...]). These modifications preserve model response content to the maximum extent while ensuring ground truth format compliance.
B.4 Entropy Analysis
Setup.
In entropy analysis, we configure the models with a sampling temperature of $0.6$, a top_ptop\_p value of $0.95$, and a maximum response length of $32768$ tokens to balance response diversity and quality. Each model generates 32 completions per problem following the avg@32 evaluation protocol, and all reported metrics (accuracy, response length, token-level entropy, and answer-level entropy) are averaged across these 32 completions and across all test problems in each benchmark.
Models.
We evaluate a diverse set of reasoning models to understand the entropy characteristics across different training paradigms and parameter scales, as summarized in the following table.
Table 11: Models evaluated in the entropy analysis.
Name Full Model Name Type Parameters
DeepSeek-1.5B DeepSeek-R1-Distill-Qwen-1.5B Base 1.5B
ProRL-1.5B Nemotron-Research-Reasoning-Qwen-1.5B RLVR 1.5B
DeepSeek-7B DeepSeek-R1-Distill-Qwen-7B Base 7B
AceReason-7B AceReason-Nemotron-7B RLVR 7B
Skywork-OR1-7B Skywork-OR1-7B RLVR 7B
DeepSeek-14B DeepSeek-R1-Distill-Qwen-14B Base 14B
AceReason-14B AceReason-Nemotron-14B RLVR 14B
Qwen2.5-32B Qwen2.5-32B Base 32B
DAPO-32B DAPO-Qwen-32B RLVR 32B
Entropy Computation.
For token-level entropy computation, we employ teacher-forcing to obtain probability estimates. Specifically, after generating the 32 completions with the specified sampling parameters, we feed each generated sequence back to the model and perform a single forward pass to compute the probability distribution over the vocabulary at each token position. Answer-level entropy is computed by first extracting the final answer from each completion using Lighteval (Habib et al., 2023), then calculating the entropy over the distribution of unique answers across the 32 completions. This approach allows us to compute both token-level and answer-level entropy directly from the model’s probability distributions without introducing additional sampling variance.
Appendix C Theoretical Limits of RLVR
C.1 Support Preservation: Why RLVR Rarely Discovers New Modes
We begin with a core limitation of RLVR: it is inherently constrained to operate within the support of the base model’s distribution. Since RLVR relies on gradient signals derived from samples generated by the base model, it cannot assign a nonzero probability to any solution that can never be sampled from $q(\cdot\mid x)$. As a result, any correct output $y^{*}$ with $q(y^{*}\mid x)=0$ remains inaccessible to policy gradient updates, regardless of reward. We formalize this intuition with Theorem C.1, which makes precise how RLVR’s reliance on the base model’s sampling prevents discovering truly new solutions.
Theorem C.1 (Support Preservation under RLVR).
Let $\pi_{\theta}(y\mid x)$ be the RLVR-trained distribution obtained via standard on-policy gradient updates on verifiable rewards $R$. Then for all $x\in\mathcal{X}$,
$\mathrm{supp}(\pi_{\theta}(\cdot\mid x))\subseteq\mathrm{supp}(q(\cdot\mid x)).$
In particular, if $q(y^{*}\mid x)=0$ for some correct solution $y^{*}$, then RLVR cannot discover $y^{*}$.
Proof.
By construction we initialize the RLVR policy to the base model as $\pi_{\theta_{0}}(y\mid x)\;=\;q(y\mid x)$. Hence
$\mathrm{supp}\bigl(\pi_{\theta_{0}}(\cdot\mid x)\bigr)=\mathrm{supp}\bigl(q(\cdot\mid x)\bigr).$
Inductive step. Assume that at some iteration $\theta$ we have
$\pi_{\theta}(y^{*}\mid x)\;=\;0\quad\text{for a particular }y^{*}.$
All standard policy‐gradient updates (e.g. REINFORCE, PPO, GRPO) take the form
$\theta^{\prime}\;=\;\theta+\eta\,\nabla_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\Bigl[R(x,y)\;-\;\beta^{-1}\log\tfrac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\Bigr],$
where $\eta$ is the learning rate. Since the outer expectation is over $y\sim\pi_{\theta}$, any $y^{*}\in\mathcal{C}$ with $\pi_{\theta}(y^{*}\mid x)=0$ is never sampled and thus contributes no gradient component. Therefore
$\pi_{\theta^{\prime}}(y^{*}\mid x)\;=\;0,$
and the support of $\pi_{\theta^{\prime}}$ remains a subset of that of $q$.
Conclusion. By induction, none of the updates can introduce positive probability mass on any $y^{*}\in\mathcal{C}$ for which $q(y^{*}\mid x)=0$. Equivalently,
$\mathrm{supp}\bigl(\pi_{\theta}(\cdot\mid x)\bigr)\;\subseteq\;\mathrm{supp}\bigl(q(\cdot\mid x)\bigr),$
indicating that any correct solution $y^{*}$ with $q(y^{*}\mid x)=0$ remains unreachable by the RLVR policy. ∎
Corollary C.2 (Asymptotic Sampling Upper Bound).
Let $\texttt{pass@}k_{p}(x)$ be the probability that at least one out of $k$ samples $y_{i}\sim p(\cdot\mid x)$ is correct, i.e. $\texttt{pass@}k_{p}(x)=1-\bigl(\Pr_{y\sim p}[R(x,y)=0]\bigr)^{k}.$ Under the conditions of Thm. C.1 and the sampling independence, we have
$\limsup_{k\to\infty}\texttt{pass@}k_{\pi_{\theta}}(x)\;\;\leq\;\;\limsup_{k\to\infty}\texttt{pass@}k_{q}(x).$
Proof.
From Thm. C.1, support preservation implies $\mathrm{supp}(\pi_{\theta}(\cdot|x))\;\;\subseteq\;\;\mathrm{supp}(q(\cdot|x)).$ Thus, for any $y\in C$, $\pi_{\theta}(y|x)>0\;\;\Longrightarrow\;\;q(y|x)>0.$
Define the total mass on correct completions by
$\pi_{\theta}(C)\;=\;\Pr_{y\sim\pi_{\theta}}[R(x,y)=1],\quad q(C)\;=\;\Pr_{y\sim q}[R(x,y)=1].$
Here, samples are assumed independent across the different draws of LLMs; otherwise, we can only assert an upper bound using union bounds. As $k\to\infty$, the pass@k success probability becomes
$\texttt{pass@}k_{\pi_{\theta}}(x)=1-(1-\pi_{\theta}(C))^{k}\;\;\longrightarrow\;\;\begin{cases}1,&\pi_{\theta}(C)>0,\\[4.30554pt] 0,&\pi_{\theta}(C)=0,\end{cases}$
and similarly for $q$.
Because support preservation ensures that any correct completion reachable under $\pi_{\theta}$ must also be reachable under $q$,
$\pi_{\theta}(C)>0\;\;\Longrightarrow\;\;q(C)>0.$
Hence, the asymptotic success probability satisfies
$\lim_{k\to\infty}\texttt{pass@}k_{\pi_{\theta}}(x)\;\;\leq\;\;\lim_{k\to\infty}\texttt{pass@}k_{q}(x).$
Theorem C.1 and Corollary C.2 prove that RLVR optimization cannot expand the search space beyond the initial support of the base model. This limitation arises because on-policy sampling means the model updates only from what it already samples — lacking representational coverage means no gradient can ever push probability mass toward truly unseen solutions. Even when rewards provide a clear training signal, RLVR cannot access or discover solutions that the base model assigns zero probability.
This manifests as a trade-off between sharpness and diversity: RLVR can improve pass@1 by concentrating mass on known high-reward modes but tends to reduce pass@k performance for larger $k$, where broader coverage is beneficial. By contrast, the base model may occasionally sample correct answers from its long-tail distribution, giving it a statistical edge under high-$k$ evaluations (Yue et al., 2025a; Liu et al., 2025). This asymptotic upper bound captures a ceiling: no matter how many samples are drawn, the RLVR-trained model cannot exceed the base model’s pass@k in the limit.
Theorem C.3 (Empirical Support Preservation).
Assume $\epsilon$ is below the finite-sample detectability threshold used in rollouts. Then, under standard sampling and update procedures with a finite sample budget, we have
$\mathrm{supp}_{\epsilon}\!\big(\pi_{\theta}(\cdot\mid x)\big)\ \subseteq\ \mathrm{supp}_{\epsilon}\!\big(q(\cdot\mid x)\big).$
Proof.
Following Zhu et al. (2025), the total update in RLVR training decomposes into
$\nabla L_{\mathrm{total}}=\nabla L_{\mathrm{PSR}}+\nabla L_{\mathrm{NSR}},$
where PSR (positive sample reinforcement) promotes correct completions and NSR (negative sample reinforcement) demotes incorrect ones while redistributing mass proportionally to the current policy. If $y\notin\mathrm{supp}_{\epsilon}(q)$, then $q(y\mid x)\leq\epsilon$, so $y$ is not $\epsilon$-detectable under the base model and will not be sampled as a positive example. Thus $\nabla L_{\mathrm{PSR}}$ has no contribution to $y$, and its probability can only change via $\nabla L_{\mathrm{NSR}}$.
NSR gradient structure. We first analyze a single decoding position. At any position with logits $z$ and probabilities $\pi_{v}$, for a sampled wrong token $y_{t}$ and learning rate $\eta$, the NSR gradient satisfies
$\frac{\partial L_{\mathrm{NSR}}}{\partial z_{v}}\propto\begin{cases}-\pi_{y_{t}}(1-\pi_{y_{t}}),&v=y_{t},\\[2.0pt] \pi_{v}\,\pi_{y_{t}},&v\neq y_{t},\end{cases}\qquad\Delta z_{v}=\eta\!\left(-\frac{\partial L_{\mathrm{NSR}}}{\partial z_{v}}\right).$
The softmax policy update under a small NSR step $\Delta z$ has the multiplicative form
$\pi^{\prime}(v)\;=\;\frac{\pi(v)\,\exp(\Delta z_{v})}{\sum_{u}\pi(u)\,\exp(\Delta z_{u})}.$
For a correct token $a\neq y_{t}$, this gives
$\Delta z_{a}=\eta\pi(a)\pi(y_{t}),\qquad\Delta z_{y_{t}}=-\eta\pi(y_{t})(1-\pi(y_{t})),\qquad\Delta z_{u}=\eta\pi(u)\pi(y_{t})\geq 0\ \ (u\notin\{a,y_{t}\}).$
Therefore,
$\frac{\pi^{\prime}(a)}{\pi(a)}\;=\;\frac{\exp(\Delta z_{a})}{\sum_{u}\pi(u)\exp(\Delta z_{u})}\;\leq\;\frac{\exp(\eta\pi(a)\pi(y_{t}))}{1-\eta\pi(y_{t})^{2}}.$
Using $\exp(\eta\pi(a)\pi(y_{t}))\leq\exp(\eta\pi(y_{t}))$ and $1/(1-x)\leq e^{2x}$ for $x\in[0,1/2]$, we obtain
$\frac{\pi^{\prime}(a)}{\pi(a)}\;\leq\;\exp\!\big(2\eta\pi(y_{t})\big).$
Iterating for $K$ steps yields the token-level bound
$\pi^{(K)}(a\mid x,y_{<t})\;\leq\;\pi^{(0)}(a\mid x,y_{<t})\,\exp\!\Big(2\eta\sum_{k=0}^{K-1}\pi^{(k)}(y_{t}\mid x,y_{<t})\Big)\;\leq\;\pi^{(0)}(a\mid x,y_{<t})\,e^{2\eta K}.$
Extension to sequences. For a full sequence $y^{\star}=(a_{1},\dots,a_{T})$, the autoregressive factorization gives
$\pi^{(K)}(y^{\star}\mid x)\;=\;\prod_{t=1}^{T}\pi^{(K)}(a_{t}\mid x,a_{<t}).$
Applying the token-level bound at each position $t$,
$\pi^{(K)}(a_{t}\mid x,a_{<t})\;\leq\;\pi^{(0)}(a_{t}\mid x,a_{<t})\,e^{2\eta K}.$
Multiplying across all $T$ positions yields
$\pi^{(K)}(y^{\star}\mid x)\;\leq\;\pi^{(0)}(y^{\star}\mid x)\,\exp(2\eta TK).$
Conclusion. Thus, if a sequence $y$ lies outside $\mathrm{supp}_{\epsilon}(q)$ so that $\pi^{(0)}(y\mid x)\leq\epsilon$, then even after $K$ NSR updates we have $\pi^{(K)}(y\mid x)\leq\epsilon e^{2\eta TK}$. As $\epsilon\rightarrow 0$, multiplying it by any finite constant still yields a vanishingly small quantity; thus, any finite multiple of $\epsilon$ is statistically negligible (i.e., undetectable in practice). Therefore, $\epsilon e^{2\eta TK}$ remains negligible for any finite $K$ and $T$, and
$\mathrm{supp}_{\epsilon}(\pi_{\theta})\;\subseteq\;\mathrm{supp}_{\epsilon}(q).$
In this sense, RLVR inherits both the inductive biases and structural limitations of its initialization. Without deliberate intervention or scaling, it remains confined to the functional expressivity of the base model. Our framework formalizes why RLVR often improves sampling efficiency but rarely produces qualitatively new reasoning capabilities.
C.2 A Variational and Support-bounded Policy Update
We now present a unified view of the RLVR objective through the lens of variational inference. This reveals why RLVR is inherently support-bounded: it makes minimal updates to the base distribution while ensuring improved performance.
Proposition C.4 (KL Projection onto Reward-Consistent Distributions).
Let $\Delta(\mathcal{Y})$ be the probability simplex over the finite output space $\mathcal{Y}$. Define the set of feasible policies that achieve at least a target expected reward $\rho$:
$\mathcal{P}_{\rho}:=\left\{p(y\mid x)\in\Delta(\mathcal{Y})\ \middle|\ \mathbb{E}_{p}[R(x,y)]\geq\rho\right\}.$
Then the solution to the variational problem, $\min_{\pi\in\mathcal{P}_{\rho}}\mathrm{KL}(\pi\,\|\,q)$, is the distribution within $\mathcal{P}_{\rho}$ that is closest in KL divergence to the base model. The optimal policy takes the form:
$\pi^{*}(y\mid x)\propto q(y\mid x)\cdot\exp(\beta R(x,y)),$
where $\beta\in\mathbb{R}_{\geq 0}$ is the dual variable associated with the reward constraint and $\beta=0$ degenerates to the base policy $q$.
Proof.
We provide two closely related derivations to illuminate the same optimal solution from both a hard-constrained and a soft-regularized perspective.
Convexity of Feasible Set $P_{\rho}$.
We first prove the convexity of $P_{\rho}$. Recall $P_{\rho}=\left\{p\in\Delta(\mathcal{Y}):\sum_{y}p(y)R(x,y)\geq\rho\right\}$, where $\Delta(\mathcal{Y})$ denotes the probability simplex over $\mathcal{Y}$.
Take any two distributions $p_{1},p_{2}\in P_{\rho}$ and let $\lambda\in[0,1]$. Consider the convex combination
$p_{\lambda}:=\lambda p_{1}+(1-\lambda)p_{2}.$
Since $\Delta(\mathcal{Y})$ is convex, we have $p_{\lambda}\in\Delta(\mathcal{Y})$.
Next, because $p_{1},p_{2}\in P_{\rho}$, it follows that
$\sum_{y}p_{1}(y)R(x,y)\geq\rho\quad\text{and}\quad\sum_{y}p_{2}(y)R(x,y)\geq\rho.$
Thus,
$\sum_{y}p_{\lambda}(y)R(x,y)=\lambda\sum_{y}p_{1}(y)R(x,y)+(1-\lambda)\sum_{y}p_{2}(y)R(x,y)\geq\lambda\rho+(1-\lambda)\rho=\rho.$
Hence $p_{\lambda}\in P_{\rho}$. This shows that $P_{\rho}$ is convex.
Convexity, existence, and strong duality.
We then verify the foundational properties of the optimization problem. Recall we wish to solve
$\min_{\pi\in P_{\rho}}\mathrm{KL}(\pi\|q),\quad\text{where }P_{\rho}=\left\{\pi\in\Delta(\mathcal{Y}):\sum_{y}\pi(y)R(x,y)\geq\rho\right\}.$
The objective function $\mathrm{KL}(\pi\|q)$ is convex in $\pi$ over the probability simplex $\Delta(\mathcal{Y})$, since relative entropy is jointly convex and thus convex in $\pi$ for fixed $q$. The feasible set $P_{\rho}$ is also convex.
Moreover, if there exists a strictly feasible distribution $\pi$ such that $\sum_{y}\pi(y)R(x,y)>\rho$, then by Slater’s condition, strong duality holds. This guarantees that the optimal value of the primal problem equals the optimal value of its Lagrangian dual, and the Karush-Kuhn-Tucker (KKT) conditions characterize the optimal solution. In typical applications—where $q$ arises from softmax-based models with full support—such strictly feasible distributions exist, ensuring that our subsequent Lagrangian approach is valid.
1) Hard-constrained formulation (projection perspective).
Consider the optimization problem:
$\min_{\pi}\mathrm{KL}(\pi\|q)\quad\text{s.t.}\quad\mathbb{E}_{\pi}[R(x,y)]\geq\rho,\quad\sum_{y}\pi(y\mid x)=1,\quad\pi(y\mid x)\geq 0.$
Using the method of Lagrange multipliers, the Lagrangian is:
$\mathcal{L}(\pi,\beta,\lambda)=\sum_{y}\pi(y\mid x)\log\frac{\pi(y\mid x)}{q(y\mid x)}-\beta\left(\sum_{y}\pi(y\mid x)R(x,y)-\rho\right)+\lambda\left(\sum_{y}\pi(y\mid x)-1\right).$
Here, we compute the derivative concerning $\pi(y\mid x)$ for fixed multipliers, thereby finding the stationary points of the Lagrangian. Specifically, we take derivative with respect to $\pi(y\mid x)$ and set it to zero:
$\log\frac{\pi(y\mid x)}{q(y\mid x)}+1-\beta R(x,y)+\lambda=0.$
Solving for $\pi$ yields:
$\pi(y\mid x)\propto q(y\mid x)\cdot\exp(\beta R(x,y)).$
2) Soft-regularized formulation (dual perspective).
Alternatively, assume RLVR solves the entropy-regularized objective
$\pi_{\theta}=\arg\max_{\pi\ll q}\mathbb{E}_{y\sim\pi}[R(x,y)]-\beta^{-1}\mathrm{KL}(\pi\,\|\,q),$
for some inverse temperature parameter $\beta>0$. Here, the constraint $\pi\ll q$ denotes that $\pi$ is absolutely continuous with respect to $q$, meaning $\pi(y\mid x)>0$ only if $q(y\mid x)>0$. [^1] Formally, absolute continuity $\pi\ll q$ ensures that the KL divergence $\mathrm{KL}(\pi\,\|\,q)$ is finite. If $\pi$ assigns positive mass to any output that $q$ assigns zero probability, the divergence becomes infinite. This condition also enforces support preservation: $\mathrm{supp}(\pi)\subseteq\mathrm{supp}(q)$.The objective is equivalent to the following minimization:
$\pi_{\theta}=\arg\min_{\pi\in\Delta(\mathcal{Y})}\mathrm{KL}(\pi\,\|\,q)-\beta\,\mathbb{E}_{y\sim\pi}[R(x,y)].$
The Lagrangian becomes
$\mathcal{L}(\pi,\lambda)=\sum_{y\in\mathcal{Y}}\pi(y)\log\frac{\pi(y)}{q(y)}-\beta\sum_{y\in\mathcal{Y}}\pi(y)R(x,y)+\lambda\left(\sum_{y\in\mathcal{Y}}\pi(y)-1\right),$
where $\lambda\in\mathbb{R}$ is the Lagrange multiplier enforcing the normalization constraint.
Taking the derivative with respect to $\pi(y)$ and setting it to zero:
$\frac{\partial\mathcal{L}}{\partial\pi(y)}=\log\frac{\pi(y)}{q(y)}+1-\beta R(x,y)+\lambda=0.$
Solving for $\pi(y)$ gives:
$\pi(y)=q(y)\cdot\exp\left(\beta R(x,y)-\lambda-1\right).$
Letting the normalization constant be:
$Z=\sum_{y^{\prime}\in\mathcal{Y}}q(y^{\prime})\cdot\exp(\beta R(x,y^{\prime})),$
we absorb constants into $Z$ and write:
$\pi_{\theta}(y\mid x)=\frac{q(y\mid x)\cdot\exp(\beta R(x,y))}{Z}.$
Both derivations recover the same exponentially tilted distribution that emphasizes high-reward completions relative to the base model. In the hard-constrained view, $\beta$ is a Lagrange multiplier tuned to meet the target reward $\rho$; in the soft-regularized view, $\beta$ sets the strength of the trade-off between reward and divergence. This completes the constructive proof of Prop. C.4.
Notably, by standard convex duality, this solution also arises as the optimizer of the entropy-regularized problem $\max_{\pi\ll q}\;\mathbb{E}_{\pi}\bigl[R(x,y)\bigr]-\frac{1}{\beta}\,\mathrm{KL}\bigl(\pi\,\|\,q\bigr)$, which softens the constraint into a penalty. Thus, RLVR can be interpreted either as a hard projection onto the closest distribution achieving the reward target, or as a soft trade-off that balances expected reward with closeness to the base model. Similar exponential tilting policy improvement oracles have been analyzed in the context of KL-regularized contextual bandits and RLHF (Zhao et al., 2024), though their focus is on sample complexity under coverage.
KL-Free Limit.
A relevant special case is the KL-free limit, where explicit KL regularization is removed ($\beta\rightarrow\infty$(Wei et al., 2023; Yu et al., 2025; Luo et al., 2025a; Yue et al., 2025b). In this regime, RLVR simplifies to a hard-filtered projection onto reward-maximizing completions.
Corollary C.5 (KL‐Free Projection).
In the limit $\beta\to\infty$, the RLVR update converges to the renormalized restriction of the base model to the correct completion set:
$\lim_{\beta\to\infty}\pi_{\beta}(y\mid x)\;=\;\frac{q(y\mid x)\,\mathbf{1}\{y\in\mathcal{C}\}}{\sum_{y^{\prime}\in\mathcal{C}}q(y^{\prime}\mid x)}.$
Proof.
Since $R(x,y)\in\{0,1\}$, we have
$\exp\bigl(\beta R(x,y)\bigr)=\begin{cases}e^{\beta}&\text{if }R(x,y)=1,\\[4.30554pt] 1&\text{if }R(x,y)=0.\end{cases}$
Thus, the RLVR distribution becomes
$\pi_{\beta}(y\mid x)=\frac{q(y\mid x)\,\exp\bigl(\beta R(x,y)\bigr)}{Z_{\beta}(x)}ni=\frac{q(y\mid x)\,\bigl[e^{\beta}\mathbf{1}\{R(x,y)=1\}+\mathbf{1}\{R(x,y)=0\}\bigr]}{Z_{\beta}(x)},$
where
$Z_{\beta}(x)=e^{\beta}\sum_{y^{\prime}:R(x,y^{\prime})=1}q(y^{\prime}\mid x)+\sum_{y^{\prime}:R(x,y^{\prime})=0}q(y^{\prime}\mid x).$
As $\beta\to\infty$, the term with $e^{\beta}$ dominates whenever there exists at least one $y$ with $R(x,y)=1$. Thus
$Z_{\beta}(x)\approx e^{\beta}\sum_{y^{\prime}\in\mathcal{C}}q(y^{\prime}\mid x).$
Similarly, in the numerator, we have
$q(y\mid x)\,\exp\bigl(\beta R(x,y)\bigr)=\begin{cases}q(y\mid x)\,e^{\beta}&\text{if }y\in\mathcal{C},\\[4.30554pt] q(y\mid x)&\text{otherwise}.\end{cases}$
Dividing by $Z_{\beta}(x)$ and taking $\beta\to\infty$, the probabilities assigned to $y$ with $R(x,y)=0$ vanish:
$\pi_{\beta}(y\mid x)\approx\begin{cases}\dfrac{q(y\mid x)\,e^{\beta}}{e^{\beta}\sum_{y^{\prime}\in\mathcal{C}}q(y^{\prime}\mid x)}=\dfrac{q(y\mid x)}{\sum_{y^{\prime}\in\mathcal{C}}q(y^{\prime}\mid x)}&\text{if }y\in\mathcal{C},\\[12.91663pt] 0&\text{otherwise}.\end{cases}$
Thus we obtain
$\lim_{\beta\to\infty}\pi_{\beta}(y\mid x)=\frac{q(y\mid x)\,\mathbf{1}\{y\in\mathcal{C}\}}{\sum_{y^{\prime}\in\mathcal{C}}q(y^{\prime}\mid x)},$
Together, Prop. C.4 and Cor. C.5 illustrate a continuum of RLVR behaviors—from softly regularized reweighting (small $\beta$) to sharply constrained filtering (large $\beta$). Even in the KL-free limit, updates remain fundamentally anchored to the base model’s distribution, preserving relative probabilities within the reward-consistent subset. Consequently, while this projection ensures stable, efficient updates, it inherently limits RLVR’s exploratory capacity. As established in Thm. C.1, RLVR remains confined to the initial support of the base model unless explicit mechanisms introduce meaningful probability mass to new regions. Thus, the variational interpretation clarifies RLVR’s strengths in improving precision and efficiency within existing capabilities, alongside its limitations in fundamentally expanding model reasoning.
C.3 Entropy–Reward Trade-off: Precision at the Cost of Answer Diversity
Another structural property of RLVR is its tendency to systematically reduce the entropy of the answer distribution. This behavior arises naturally from reward optimization, which statistically favors sharper distributions concentrated on high-reward completions. While such entropy reduction is beneficial in domains like board games or math—where precision is paramount—it may also suppress valuable diversity in contexts that benefit from broader coverage or multiple valid outputs, such as story or dialogue generation (Chen et al., 2023) and coding copilots (Peng et al., 2023).
Theorem C.6 (Entropy Reduction and Precision–Coverage Trade-off).
Assume a finite output space $\mathcal{Y}$ and define the Shannon entropy of a distribution as $\mathcal{H}[p]:=-\sum_{y\in\mathcal{Y}}p(y\mid x)\log p(y\mid x).$ Then the following statements hold:
  1. (a)
    Entropy reduction. Any RLVR update $\pi_{\theta}$ satisfies
    $\mathcal{H}[\pi_{\theta}]\;\leq\;\mathcal{H}[q],$
    with equality only if the reward is constant on the support of $q$.
  2. (b)
    Trade-off with coverage. Lower entropy increases sampling precision for small budgets, but for large $k$, reduces the diversity of explored outputs—potentially missing alternative correct completions.
Proof.
(a) Entropy reduction. Consider the exponentially tilted distribution
$\pi_{\theta}(y\mid x)=\frac{q(y\mid x)\exp(\beta R(x,y))}{Z},\quad\text{with}\quad Z=\sum_{y\in\mathcal{Y}}q(y\mid x)\exp(\beta R(x,y)).$
By standard properties of KL divergence,
$\mathrm{KL}(\pi_{\theta}\|q)=\sum_{y}\pi_{\theta}(y\mid x)\log\frac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\;\;\geq\;\;0.$
Rearranging gives
$\mathcal{H}[\pi_{\theta}]=\mathcal{H}[q]-\mathrm{KL}(\pi_{\theta}\|q)\;\;\leq\;\;\mathcal{H}[q].$
Thus, any such RLVR update decreases entropy relative to the base distribution, unless the reward is constant (in which case $\pi_{\theta}=q$).
(b) Trade-off with diversity at different sampling budgets. The RLVR-trained policy sharpens the probability mass around high-reward completions. Explicitly,
$\pi_{\theta}(y\mid x)\;\propto\;q(y\mid x)\exp(\beta R(x,y)),$
where $\beta>0$ controls concentration.
  • Small sampling budgets ($k=1$): The increased probability on high-reward outputs generally improves single-shot success rates. Formally,
    $\texttt{pass@1}_{\pi_{\theta}}(x)=\sum_{y:R(x,y)=1}\pi_{\theta}(y\mid x)\;\;>\;\;\sum_{y:R(x,y)=1}q(y\mid x)=\texttt{pass@1}_{q}(x),$
    provided the reweighting boosts correct completions relative to incorrect ones.
  • Large sampling budgets ($k\gg 1$): However, reduced entropy leads to concentration on fewer modes. As $\beta$ grows, $\pi_{\theta}$ may collapse onto a narrow subset of correct completions, neglecting other valid solutions accessible under the more dispersed $q$. Thus,
    $\limsup_{k\to\infty}\texttt{pass@}k_{\pi_{\theta}}(x)<\limsup_{k\to\infty}\texttt{pass@}k_{q}(x),$
    under typical conditions of entropy reduction and selective mass shifting.
  • Loss of tail coverage: In particular, if there exist rare but correct completions that have small mass under $q$ but are further downweighted (or eliminated) by the tilting, then the total mass on correct completions can decrease:
    $\pi_{\theta}(C)<q(C),\quad C=\{y:R(x,y)=1\}.$
    This restricts the long-run probability of recovering diverse solutions via large $k$ sampling.
Conclusion.
This establishes a trade-off: RLVR improves sampling efficiency by concentrating probability on high-reward outputs (increasing pass@1), but this comes at the cost of reduced entropy and narrower exploration of the solution space (potentially lowering pass@k for large $k$). Empirical studies confirm this phenomenon in settings like code generation and symbolic reasoning, where many semantically distinct correct completions exist. ∎
This trade-off underpins RLVR’s empirical strengths in tasks with narrowly defined optimal solutions, such as mathematical proofs or tactical game endgames (where precision is paramount), while also emphasizing the need for explicit diversity mechanisms in more open-ended domains, such as code generation, creative writing (Feizi et al., 2023; Ding et al., 2024), or brainstorming (Chang & Li, 2025). Importantly, entropy reduction is not inherently undesirable: when a task admits a unique correct solution, lower answer-level entropy simply reflects desirable convergence. Importantly, even in multi-solution domains, concentrating mass on a narrower set may still be desirable under constrained compute budgets. However, our results show that entropy reduction can still lead to empirical support shrinkage even in predominantly single-solution domains like math, where RLVR sometimes fails to recover valid completions still accessible to the more diverse base model. This highlights that entropy-induced narrowing is a general phenomenon, not limited to multi-solution tasks, underscoring the broader need for explicit exploration or diversity-promoting strategies.
C.4 Estimating the Sampling Threshold $\epsilon$ from pass@k
We provide a statistical analysis of the threshold $\epsilon$ in the pass@k sampling. Suppose we sample $k$ times from a model $\pi(\cdot\mid x)$, and let $y^{*}\in\mathcal{C}$ be a correct completion with unknown probability $p=\pi(y^{*}\mid x)$. If $y^{*}$ is not observed in any of those $k$ samples, we can upper bound $p$ using the following argument.
The probability of not sampling $y^{*}$ in a single trial is 1p1-p, so the probability of missing it in all $k$ independent trials is $(1-p)^{k}$. To ensure this event occurs with probability at most $\zeta$, we solve:
$(1-p)^{k}\leq\zeta.$
Taking logarithms of both sides:
$k\cdot\log(1-p)\leq\log\zeta.$
Using the inequality $\log(1-p)\leq-p$ for $p\in(0,1)$, we get:
$k\cdot(-p)\geq\log\zeta\quad\Rightarrow\quad p\leq\frac{-\log\zeta}{k}.$
Consequently, if the correct completion $y^{*}$ is not observed in $k$ samples, then with confidence $1-\zeta$, its probability satisfies:
$\boxed{\pi(y^{*}\mid x)\leq\frac{-\log\zeta}{k}.}$
Example.
If $k=8192$ in the math reasoning tasks and we desire 95% confidence (i.e., $\zeta=0.05$), then
$\pi(y^{*}\mid x)\leq\frac{-\log(0.05)}{8192}\approx\frac{2.996}{8192}\approx 3.66\times 10^{-4}.$
Appendix D The use of LLMs
This study utilized large language models solely for providing assistance with minor language enhancements. All content has undergone human review, verification, and further modification.